OpenStudy (anonymous):

Explain, in complete sentences, how you would factor 400x8 - 36.

OpenStudy (anonymous):

\[4\left[ \left( 10 x^4 \right)^2-3^2 \right]=4\left( 10 x^4+3 \right)\left( 10x^4-3 \right)\] \[=4\left[ \left( 10x^4 \right)+3 \right]\left[ \left( \sqrt{10} x^2\right)^2-\left( \sqrt{3} \right)^2 \right]\] =?

OpenStudy (triciaal):

is that x to the 8th ?

OpenStudy (triciaal):

find factors of each term 400x^8 = (2*10*x^4)^2 and 36 = 6^2 difference of 2 squares a^2 -b^2 = (a + b)(a - b) a = 20x^4 and b = 6